The cooling system in a truck contains 4 gallons of coolant that is 30% antifreeze. How much must be withdrawn and replace with 100% antifreeze to bring the coolant in the system to 50% antifreeze?
I have no idea where to start.

Answer by lwsshak3(11628)

The cooling system in a truck contains 4 gallons of coolant that is 30% antifreeze. How much must be withdrawn and replace with 100% antifreeze to bring the coolant in the system to 50% antifreeze?
let x=gallons of 30% coolant to be removed and replaced with 100% antifreeze.
then 4-x gallons of 30% coolant remains initially
then adding x gallons of 100% antifreeze will give a 50% antifreeze coolant
.3(4-x)+100%x=.5*4
1.2-.3x+x=2
.7x=.8
x=8/7 gal
ans: 8/7 gal of 30% antifreeze coolant must be removed and replaced with 100% antifreeze to bring the coolant in the system to 50% antifreeze?